OpenStudy (anonymous):

Mrs. Computer is 3 times older than her daughter, Mousy. The sum of their ages is 52. How old is Mousy?

OpenStudy (anonymous):

mousy is 13 yrs old

OpenStudy (mathteacher1729):

COMPUTER's age = three times MOUSY's age as an equation can be written as: C = 3 * M "Sum of their ages" means "when you add their ages together". COMPUTER's age + MOUSY's age = 52 as an equation that is C + M = 52 So now you have two equations: C = 3m C + M = 52 Can you solve it from here?